Two cars start from the same point at the same time and travel in opposite directions. The faster car travels 12 kilometers per hour faster than the other, if after 6 hours the cars are 648 kilometers apart, what is the speed of each car?
Answer by rfer(16322)

648/6=108
x+x+12=108
2x=108-12
2x=96
x=96/2
x=48 kph
x+12=60 kph